OpenStudy (anonymous):

Convert (-4,6) and (-9,6) to polar points

OpenStudy (anonymous):

@Koikkara

OpenStudy (koikkara):

@aravin.....hope u know this bttr than me... http://facultypages.morris.umn.edu/~mcquarrb/teachingarchive/Precalculus/Lectures/PolarCoordinates.pdf refer....

OpenStudy (anonymous):

i did r=sqrtx^2+y^2 but it didn't work

OpenStudy (anonymous):

What didn't work?

OpenStudy (anonymous):

\[\sqrt{(x^2)+(y^2)}\]

OpenStudy (zehanz):

Hey, @abc123! Just do it like the other one! r = sqrt(x²+y²)=sqrt((-4)²+6²) Maybe problem w. brackets ?

OpenStudy (anonymous):

thats what i did. i got 2sqrt13 but it said i was wrong

OpenStudy (zehanz):

Well, \(\sqrt{52}=\sqrt{4 \cdot 13}=\sqrt{4}\cdot\sqrt{13}=2\sqrt{13}\), so that is OK!

OpenStudy (anonymous):

but it says i am wrong so i don't know what to do

OpenStudy (zehanz):

What angle did you get? Could that be the problem?

OpenStudy (anonymous):

arccos(-4/2sqrt13)

OpenStudy (zehanz):

That is about 2.16 radians. That is also OK. You just gave the right answer. Maybe the error occured after you gave the conversion of the other point, (-9, 6)?

OpenStudy (anonymous):

no its two seperate questions

OpenStudy (koikkara):

@ZeHanz ...i applaud You for the Explanation....Keep it up...!

OpenStudy (zehanz):

Well @abc123!, sometimes you are right and the book, teacher, or whatever computer program is wrong! Don't bother with it anymore, and go on with what you were doing!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i am also having probs with (13,6) and (20,2)

OpenStudy (zehanz):

@Koikkara: thx for the support!

OpenStudy (zehanz):

@abc123! What r's have you found?

OpenStudy (anonymous):

for (13,6) i have sqrt205 and for (20,2) i have 2sqrt101. i know they r right. just the angles r wrong

OpenStudy (zehanz):

OK, let's look into that. What values did you get?

OpenStudy (anonymous):

arccos(13/sqrt205) and arccos(20/2sqrt101)

OpenStudy (zehanz):

There could be a slight problem with the last one: type arccos(20/(2sqrt101)) to be sure 20 gets divided not only by 2, but also by sqrt101.

OpenStudy (anonymous):

no i converted it to a decimal

OpenStudy (zehanz):

and got .996686525?

OpenStudy (anonymous):

.09966865.....

OpenStudy (zehanz):

Right again. Are doing some kind of online test, where you have to type in your answers?

OpenStudy (anonymous):

its online hw not a test

OpenStudy (zehanz):

Are you shown the "right" answers at any time?

OpenStudy (anonymous):

after the session is closed which is tuesday at 10

OpenStudy (zehanz):

And you have to give the answers in decimal format? How many decimals?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

it wont let me put in arccos so i think it has to be a decimal

OpenStudy (zehanz):

Could you try arctan? I.e. arctan(6/13) for the point (13,6)

OpenStudy (anonymous):

it says that 'Can't take acos of 100.498756211209" for arccos(20/2sqrt101) but i got the angle for (13,6)

OpenStudy (anonymous):

arccos worked for (13,6)

OpenStudy (zehanz):

We've found it!

OpenStudy (anonymous):

@abc123! To Convert from Cartesian to Polar When you know a point in Cartesian Coordinates (x,y) and want it in Polar Coordinates (r,θ) you solve a right triangle with two known sides. Example: What is (12,5) in Polar Coordinates? Use Pythagoras Theorem to find the long side (the hypotenuse): r2 = 122 + 52 r = √ (122 + 52) r = √ (144 + 25) = √ (169) = 13 Use the Tangent Function to find the angle: tan( θ ) = 5 / 12 θ = tan-1 ( 5 / 12 ) = 22.6° (to one decimal) What is tan-1 ? It is the Inverse Tangent Function. Tangent takes an angle and gives you a ratio, Inverse Tangent takes a ratio (like "5/12") and gives you an angle. Answer: the point (12,5) is (13, 22.6°) in Polar Coordinates. So to convert from Cartesian Coordinates (x,y) to Polar Coordinates (r,θ): r = √ ( x2 + y2 ) θ = tan-1 ( y / x )

OpenStudy (zehanz):

It's what I told you earlier about 20/2sqrt101. it should be: 20/(2sqrt101)

OpenStudy (anonymous):

ok. let me try

OpenStudy (zehanz):

@niksva: thanks for the help, but we've got a "technical" problem here, I'm afraid...

OpenStudy (anonymous):

yup it worked. but now i can't get anything for (-9,6) and (-4,6)

OpenStudy (anonymous):

i got 2sqrt13 for (-4,6) and 3sqrt13 for (-9,6). it says its wrong

OpenStudy (zehanz):

The r's we can do w/o calculator: sqrt(81+36)=sqrt(117)=3sqrt(13) sqrt(16+36)=sqrt(52)=2sqrt(13)

OpenStudy (anonymous):

yeah thats what i got

OpenStudy (zehanz):

OK now the thetas: arctan(6/(-9))=-.5880. Adding pi gives 2.5536 arctan(6/(-4))=-.9828. Adding pi gives 2.1588

OpenStudy (anonymous):

how do i put that in without a decimal

OpenStudy (zehanz):

arctan(6/(-9))+pi ?

OpenStudy (zehanz):

Can you input a greek letter pi?

OpenStudy (anonymous):

yeah i can and its wrong

OpenStudy (zehanz):

No, it's right, but the stupid prog doesn't know...

OpenStudy (anonymous):

ok whatever thanks for helping

OpenStudy (zehanz):

YW!

OpenStudy (phi):

try putting in parens i got 2sqrt13 for (-4,6) try 2 sqrt(13) for the angle try acos(-4/sqrt(52))

OpenStudy (anonymous):

ok